OpenStudy (anonymous):

what is the exact value of cos 54* cos 8* + sin 53 * sin 8* *= Degrees

OpenStudy (anonymous):

That looks a lot like a trig identity.. \(cos(\theta - \phi)\) probably.

OpenStudy (anonymous):

Well where doing functions of the sum of two angles... if that means anything Trigonnometric applications. of some sort and i know the chart is involved with the 30 45 60 sin , cos, tan kinda thing what is the thing next to beta.?

OpenStudy (anonymous):

Yeah, if you look in your book, or your notes you should see a formula that looks like this one except without the numbers. The formula you have here is the one for the sum/difference of two angles for cosine. \[cos(\theta \pm \phi) = cos(\theta)cos(\phi) \mp sin(\theta)sin(\phi)\] In this case, \(\theta=53\) and \(\phi=8\). Since you are adding the product of the sines to the product of the cosines you just want to take the cosine of the difference between the two. 53-8 = 45. So the answer is the cos of 45 degrees.